# Generalized Pulse Design in Floquet Engineering: Application to Interacting Spin Systems

### v1

@@ III.1.3 Time-Ordering of Global Pulses
[…]
@@ III.1.4 Convergence Considerations
The general problem of determing the radius of convergence of high frequency expansions at infinite times is non-trivial and remains an open question (See Refs. 29, 23, 24, 41 for discussions). However, high frequency expansions can be safely applied in practice for sufficiently short total time scales $NT$ such that issues related to the radius of convergence [42] are avoided. In the following we therefore assume that, in practice, a quantum simulator evolves for a finite amount of time such that perturbative truncations discussed below are small and the system does not reach a long time asymptotic regime [24, 41].
@@ III.2 Effective Hamiltonian from Strong Driving
This section uses the above assumptions to construct a formalism for pulse engineering effective Hamiltonians starting with Eq. (1). We require the potential to be cyclic and periodic as discussed in Sec. III.1.1. The perturbing potential $\hat{V}(t)$ is assumed to have the form given in Sec. III.1.2, i.e., we assume Eqs. (12)-(13). We proceed via a gauge transformation on the Schrödinger equation defined by Eqs. (3). We expand $\hat{H}_{\text{eff}}$ and $\hat{K}(t)$ in powers of inverse frequency following our assumption in Sec. III.1.4:
@@
By using identities of Lie algebras and equating powers of $\omega^{-1}$, $\hat{K}^{(0)}$ must solve the (truncated) lowest order differential equation (see Appendix B for details):
$\frac{\partial\hat{K}^{(0)}}{\partial t}=\omega\sum_{j}\sum_{\alpha\in\mathcal{A}}g^{\alpha}_{j}(t)\hat{\sigma}^{(\alpha)}_{j},$ (17)
which implies that the kick operator has the form:
$\hat{K}^{(0)}$ $=\sum_{j}\sum_{\alpha\in\mathcal{A}}G^{\alpha}_{j}(t)\hat{\sigma}^{(\alpha)}_{j},$ (18)
with:
[…]
such that these $G^{\alpha}_{j}(t)$ are also periodic and cyclic (See Fig. 2 for an example). If we take averages and only expand to lowest order, $\omega^{0}$, we obtain terms proportional to the average value of $g^{\alpha}_{j}(t)G^{\beta}_{j^{\prime}}(t)$ and the average value of $G^{\alpha}_{j}(t)G^{\beta}_{j^{\prime}}(t)$. We can choose the $g^{\alpha}_{j}(t)$ so that
$g^{\beta}_{j^{\prime}}\cdot G^{\alpha}_{j}$ $=0$
$G^{\beta}_{j^{\prime}}\cdot G^{\alpha}_{j}$ $=(G^{\alpha}_{j})^{2}\delta_{jj^{\prime}}\delta_{\alpha\beta},$
[…]
Figure 3: A schematic depicting the expansion procedure used to derive Eqs. (18) and  (20). The first row ($n=0$) in the central rectangle shows that Eq. (20) retains all nested commutators to lowest order in inverse frequency, $\omega^{0}$. The rectangle on the right shows that the lowest order kick operator must also be kept. The remaining rows ($n>0$) are truncated in our formalism. The dashed line depicts an example $\mathcal{O}(\omega^{-1})$ correction to highlight the need for higher order expansions of the kick operator, $K^{(1)}$, when retaining $n=1$ terms.
Keeping all $\mathcal{O}(\omega^{0})$ terms in the expansion yields the following high frequency effective Hamiltonian (See Appendix B for a derivation):
$\hat{H}_{\text{eff}}=\hat{H}_{0}+\sum_{j}\sum_{\alpha\in\mathcal{A}}F^{\alpha}_{j}[\hat{H}_{0}]+\mathcal{O}(\omega^{-1}),$ (20)
@@
$F^{\alpha}_{j}[\hat{H}]=\sum_{p=1}^{\infty}\frac{(-1)^{p}}{(2p)!}\overline{\left(G^{\alpha}_{j}\right)^{2p}}[\hat{\sigma}^{(\alpha)}_{j},...[\hat{\sigma}^{(\alpha)}_{j},\hat{H}]...],$ (21)
[…]
Equations (18) and  (20) establish central results. These equations, under the assumptions discussed in Sec. III.1, are exact to $\mathcal{O}(\omega^{0})$. The green rectangular regions in Fig. 3 show the terms retained in Eqs. (18) and  (20). Fig. 3 also shows terms that need to be retained to include $\mathcal{O}(\omega^{-1})$ corrections in $\hat{H}_{\text{eff}}$. The dashed line highlights an example higher order correction to emphasize that rigorous inclusion of such terms requires solving for both $\hat{K}^{(0)}$ and $\hat{K}^{(1)}$. Appendix B also discusses the higher order $\mathcal{O}(\omega^{-1})$ corrections in the derivation. Eqs. (18) and  (20) are also very general. Specifically, Eqs. (18) and  (20) can be applied in a non-stroboscopic setting–without further assumptions regarding time scales–and to general models captured by Eq. (4).
The infinite series for $F^{\alpha}_{j}$, Eq. (21), must be evaluated to all orders to allow arbitrary pulse shapes. But an alternative simple model arises by truncating the series in $p$. Keeping only the $p=1$ term leads to:
$\hat{H}_{\text{eff}}\approx\hat{H}_{0}-\frac{1}{2}\sum_{j}\sum_{\alpha\in\mathcal{A}}\overline{(G^{\alpha}_{j})^{2}}\left[\hat{\sigma}^{(\alpha)}_{j},\left[\hat{\sigma}^{(\alpha)}_{j},\hat{H}_{0}\right]\right],$ (22)
where the corrections in the expansion are proportional to $\overline{(G^{\alpha}_{j})^{4}}$. As discussed above, pulse profile choices can implicitly introduce time scales and require retaining all $p$ but Eq. (22) will suffice if $\overline{(G^{\alpha}_{j})^{4}}$ and higher order corrections in $p$ converge rapidly. But if they do not converge rapidly because non-trivial time scales are introduced by the pulse profiles (See, e.g.,  Ref. 9), there will be significant corrections to Eq. (22) that must be accounted for. We now turn to the XXZ models where we sum the nested commutators in Eq. (21) explicitly to derive analytic control conditions for $\hat{H}_{\text{eff}}$ at all orders of $p$.
@@ IV Application to XXZ Spin Models
We now apply our formalism to XXZ spin models. We then discuss example pulses and closed form solutions that allow effective Hamiltonian engineering of Eq. (4). We focus on spatially global pulses as experimentally relvant examples.
@@ IV.1 General Conditions for Engineering
[…]
@@
which specifies the effective Hamiltonian. Lastly, we remark that Eq. (24) is broad enough in $A$ and $B$ (by inspecting their forms) that one can engineer on $\hat{H}_{XY}$ or $\hat{H}_{XXZ}$ to produce a large class of effective Hamiltonians [see also Eq. (69)]. For example, for $A=0$, we obtain the special case: $\hat{H}_{\text{eff}}^{\text{S}}\rightarrow\hat{H}_{ZZ}$. For $B=0$, we obtain $\hat{H}_{\text{eff}}^{\text{S}}\rightarrow\hat{H}_{XY}$ and for $A=B$, $\hat{H}_{\text{eff}}^{\text{S}}$ becomes the Heisenberg model: $\sim\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}[\hat{\sigma}^{(1)}_{j}\hat{\sigma}^{(1)}_{j^{\prime}}+\hat{\sigma}^{(2)}_{j}\hat{\sigma}^{(2)}_{j^{\prime}}+\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}]$.
@@ IV.2.1 Cosine Wave
$J_{0}(z)$ takes on values (approximately) in the interval $[-0.4,1]$. If we determine the set of $s$ such that $(s+3)/(s-1)$ takes on values in this interval, we can determine the set of $\hat{H}_{XXZ}$ Hamiltonians that can be engineered into $\hat{H}_{ZZ}$ using Eq. (28). Doing this, we obtain the set $s\in(-\infty,-13/7]$. Note in particular as $s\rightarrow\pm\infty$, we have $v\rightarrow 0$, which makes sense since this corresponds to $J_{\perp}\rightarrow 0$, so that $\hat{H}_{0}=\hat{H}_{ZZ}$. The restriction to $s<0$ implies that a cosine wave can produce an Ising model from the XXZ model only if there is a relative minus sign between $J_{\perp}$ and $J_{z}$.
@@
Figure 4: The solid lines plot the surfaces of solutions for $J_{z}/J_{\perp}$ obtained from Eq. (27) plotted against the pulse strength, $v$. The pulse is assumed to be a cosine shape, Eq. (28). In the top, middle, and bottom panles the dashed lines plot the truncated approximations terminating at $p_{max}=1$, $p_{max}=8$, and $p_{max}=16$, respectively. The vertical dashed lines indicate a pole in the truncated expansion. As expected, when $v\rightarrow 0$, lower order approximations are good fits to the solution surface. On the other hand, every truncated solution incorrectly predicts the existence of values of $v$ for $s>0$ (i.e., $J_{z}$ and $J_{\perp}$ having the same sign) which are pushed further and further out to $v\rightarrow\infty$. To determine all admissible values of $v$ which effectively engineer an Ising-type interaction from a given form of $H_{XXZ}$, we must keep all terms in the expansion.
@@
Appendix D briefly reviews recent quantum simulation proposals and experiments realizing XXZ models with $\mathcal{V}_{j,j^{\prime}}=|j-j^{\prime}|^{-3}$. Trapped circular Rydberg atoms can be engineered into an XXZ model where there can be a relative minus sign between $J_{\perp}$ and $J_{z}$ [43]. Polar molecules, however, defining pseudospins in the simplest setup (rotational states with total angular momenta $\mathcal{N}=0,1$ and zero $z$-component of angular momentum) do not always allow a relative minus sign between $J_{\perp}$ and $J_{z}$. We therefore find that Eq. (28) does not allow a conversion from $H_{XXZ}$ to $H_{ZZ}$ for dipolar interacting molecules. But this condition can be generically met in bipartite spin model approximations to trapped polar molecules, e.g., a nearest neighbor interaction [44] $\mathcal{V}_{j,j^{\prime}}\approx\delta_{j,j^{\prime}\pm 1}$ on a chain lattice, by performing a gauge transformation on every other site on the chain: $\sigma^{(1)}_{2j}\rightarrow-\sigma^{(1)}_{2j}$ and $\sigma^{(2)}_{2j}\rightarrow-\sigma^{(2)}_{2j}$. The gauge transformation allows a relative minus sign between $J_{\perp}$ and $J_{z}$ (on a bipartite model). We therefore conclude that for the global cosine pulse, dipolar XXZ models engineered by circular Rydberg atom arrays can be driven into an Ising regime whereas only bipartite models of interacting $\mathcal{N}=0,1$ Alkali-based polar molecules can be converted from an XXZ model to an Ising model.
@@
Figure 4 uses a strong driving cosine pulse to highlight the need to retain all orders in $p$ in Eq. (21) in a broad parameter search designed to engineeer models. One could in principle use the $p_{max}=1$ term–i.e., Eq. (22)–to attempt to engineer the XXZ model into an Ising model, but this leads to qualitatively incorrect results in a pulse parameter sweep. In particular, the first-order solution to Eq. (27) suggests a smooth set of values for $v$ to select from that effectively engineers all possible ratios of $J_{z}$ to $J_{\perp}$ up to $J_{z}=J_{\perp}$, as seen by the top panel in Fig. 4. However, even truncating at large but finite $p_{max}$ still predicts the existence of values of $v$ which could in principle engineer specific forms of the $H_{XXZ}$, but as one keeps more and more terms in Eq. (25), these solutions are pushed out to $v\rightarrow\infty$. For large $v$, the solutions are non-physical and this can only be demonstrated by keeping all orders in $p$ in wide parameter searches over all $v$.

### v2

@@ III.1.3 Form of the Kick Operator
where the pulse turns on and off the Pauli operators indexed by $\alpha$ in distinct time windows $t\in[(\alpha-1)T,\alpha T]$. We choose the $g(t)$ to satisfy the conditions: i) $\int_{(\alpha-1)T}^{\alpha T}g(t)dt=0$ which means that each $g^{\alpha}_{j}(t)$ has a subcycle time of $T$; ii) $G(t)$ has half-wave symmetry, so that $G(\frac{2\alpha+1}{2}T-\delta t)=-G(\frac{2\alpha+1}{2}T+\delta t)$, with $0<\delta t<T/2$. An example time ordering is depicted in Fig. 2 for $N=2$ and $g(t)$ is a (translated) square wave.
@@ III.1.4 Convergence Considerations
The general problem of determining the radius of convergence of high frequency expansions at infinite times is non-trivial and remains an open question (See Refs. rahav2003effective, goldman2014periodically, bukov2015universal, KUWAHARA2016 for discussions). However, high frequency expansions can be safely applied in practice for sufficiently short total time scales $NT$ such that issues related to the radius of convergence [MOAN2001a] are avoided. In the following we therefore assume that, in practice, a quantum simulator evolves for a finite amount of time such that perturbative truncations discussed below are small and the system does not reach a long time asymptotic regime [bukov2015universal, KUWAHARA2016].
@@ III.2 Effective Hamiltonian from Strong Driving
This section uses the above assumptions to construct a formalism for pulse engineering effective Hamiltonians starting with Eq. (1). We require the potential to be cyclic and periodic, as discussed in Sec. III.1.1. The perturbing potential $\hat{V}(t)$ is assumed to have the form given in Sec. III.1.2 and Sec. III.1.3, i.e., we assume Eqs. (12)-(13) and Eqs. (14)-(16). We proceed via a gauge transformation on the Schrödinger equation defined by Eqs. (3). We expand $\hat{H}_{\text{eff}}$ and $\hat{K}(t)$ in powers of inverse frequency following our assumption in Sec. III.1.4:
@@
If we employ our assumptions in Sec. III.1, take the time-average of the effective Hamiltonian, and only expand to lowest order, $\omega^{0}$, we obtain terms proportional to the average value of $g^{\alpha}_{j}(t)G^{\beta}_{j^{\prime}}(t)$ and the average value of $G^{\alpha}_{j}(t)G^{\beta}_{j^{\prime}}(t)$. We can choose the $g^{\alpha}_{j}(t)$ so that:
$\overline{g^{\beta}_{j^{\prime}}G^{\alpha}_{j}}$ $=0$
$\overline{G^{\beta}_{j^{\prime}}G^{\alpha}_{j}}$ $=(G^{\alpha}_{j})^{2}\delta_{jj^{\prime}}\delta_{\alpha\beta},$ (19)
where the overline denotes the time average value of a function: $\overline{f}=T^{-1}\int_{0}^{T}f(t)dt$, and we suppress the $t$-argument for notational convenience. For example, choosing $g_{j}^{\alpha}(t)=v\cos(\omega t)$ ($v$ a dimensionless parameter) when $t\in[(\alpha-1)T,\alpha T]$ and $g_{j}^{\alpha}(t)=0$ otherwise satisfies this condition.
Figure 3: A schematic depicting the expansion procedure used to derive Eqs. (14) and  (20). The first row ($n=0$) in the central rectangle shows that Eq. (20) retains all nested commutators to lowest order in inverse frequency, $\omega^{0}$. The rectangle on the right shows that the lowest order kick operator must also be kept. The remaining rows ($n>0$) are truncated in our formalism. The dashed line depicts an example $\mathcal{O}(\omega^{-1})$ correction to highlight the need for higher order expansions of the kick operator, $\hat{K}^{(1)}$, when retaining $n=1$ terms.
Keeping all $\mathcal{O}(\omega^{0})$ terms in the expansion yields the following time-averaged high frequency effective Hamiltonian (See Appendix B for a derivation):
$\overline{\hat{H}_{\text{eff}}}=\hat{H}_{0}+\sum_{j}\sum_{\alpha\in\mathcal{A}}F^{\alpha}_{j}[\hat{H}_{0}]+\mathcal{O}(\omega^{-1}),$ (20)
@@
$F^{\alpha}_{j}[\hat{H}]=\sum_{p=1}^{\infty}\frac{(-1)^{p}}{(2p)!}\overline{\left(G^{\alpha}_{j}\right)^{2p}}[[\hat{\sigma}^{(\alpha)}_{j},\hat{H}]]_{2p},$ (21)
where the notation $\hat{A},\hat{B}]]_{p}=[\hat{A},...[\hat{A},\hat{B}$]...] indicates a nested commutator occurring $p$ times such that:
$[[\hat{A},\hat{B}]]_{0}$ $=\hat{B}$,
$[[\hat{A},\hat{B}]]_{1}$ $=[\hat{A},\hat{B}$],
$[[\hat{A},\hat{B}]]_{2}$ $=[\hat{A},[\hat{A},\hat{B}$]],
$[[\hat{A},\hat{B}]]_{3}$ $=[\hat{A},[\hat{A},[\hat{A},\hat{B}$]]],
$\vdots$
$F^{\alpha}_{j}[\hat{H}$] enables us to think of the engineering process as a transformation on the operator algebra over the state space. Eq. (20) must also be paired with the solution to the differential equation for $\hat{K}^{(0)}$, Eq. (14), to compute observables. Note that the effective Hamiltonian applies only in the time average sense and that this approach differs from the use of term-by-term cancellation of time dependence without averaging (see, e.g., Ref. [Bukov2016]).
Equations (14) and  (20) establish central results. These equations, under the assumptions discussed in Sec. III.1, are exact to $\mathcal{O}(\omega^{0})$. The green rectangular regions in Fig. 3 show the terms retained in Eqs. (14) and  (20). Fig. 3 also shows terms that need to be retained to include $\mathcal{O}(\omega^{-1})$ corrections in $\overline{\hat{H}_{\text{eff}}}$. The dashed line highlights an example higher order correction to emphasize that rigorous inclusion of such terms requires solving for both $\hat{K}^{(0)}$ and $\hat{K}^{(1)}$. Appendix B also discusses the higher order $\mathcal{O}(\omega^{-1})$ corrections in the derivation. Eqs. (14) and  (20) are also very general. Specifically, Eqs. (14) and  (20) can be applied in a non-stroboscopic setting–without further assumptions regarding time scales–and to general spin models $H_{0}$ and, as a concrete example, models captured by Eq. (4).
The infinite series in Eq. (21) must be evaluated to all orders to allow arbitrary pulse shapes. But an alternative simple model arises by truncating the series in $p$. Keeping only the $p=1$ term leads to:
$\overline{\hat{H}_{\text{eff}}}\approx\hat{H}_{0}-\frac{1}{2}\sum_{j}\sum_{\alpha\in\mathcal{A}}\overline{(G^{\alpha}_{j})^{2}}\left[\hat{\sigma}^{(\alpha)}_{j},\left[\hat{\sigma}^{(\alpha)}_{j},\hat{H}_{0}\right]\right],$ (22)
where the corrections in the expansion are proportional to $\overline{(G^{\alpha}_{j})^{4}}$. As discussed above, pulse profile choices can implicitly introduce time scales and require retaining all $p$ but Eq. (22) will suffice if $\overline{(G^{\alpha}_{j})^{4}}$ and higher order corrections in $p$ converge rapidly. But if they do not converge rapidly because non-trivial time scales are introduced by the pulse profiles (See, e.g.,  Ref. GEIER2021c), there will be significant corrections to Eq. (22) that must be accounted for. We now turn to the XXZ models where we sum the nested commutators in Eq. (21) explicitly to derive analytic control conditions for $\overline{\hat{H}_{\text{eff}}}$ at all orders of $p$.
@@ IV Application to XXZ Spin Models
We now apply our formalism to XXZ spin models. We then discuss example pulses and closed form solutions that allow effective Hamiltonian engineering of Eq. (4). We focus on spatially global pulses as experimentally relevant examples.
@@ IV.1 General Conditions for Engineering
Here $\delta J=J_{z}-J_{\perp}$ and $G=\omega\int^{t}g(t^{\prime})dt^{\prime}$. $U$ arises as a consequence of the sum in Eq. (21) and abstains from a particular choice for $g(t)$. Implicitly, this means that the coefficients $A$ and $B$ depend on the form of the pulses chosen, which is to be expected (see Appendix C).
@@
which specifies the effective Hamiltonian. Lastly, we remark that Eq. (24) is broad enough in $A$ and $B$ (by inspecting their forms) that one can engineer on $\hat{H}_{XY}$ or $\hat{H}_{XXZ}$ to produce a large class of effective Hamiltonians [see also Eq. (74)]. For example, for $A=0$, we obtain the special case: $\overline{\hat{H}^{\text{S}}_{\text{eff}}}\rightarrow\hat{H}_{ZZ}$. For $B=0$, we obtain $\overline{\hat{H}^{\text{S}}_{\text{eff}}}\rightarrow\hat{H}_{XY}$ and for $A=B$, $\overline{\hat{H}^{\text{S}}_{\text{eff}}}$ becomes the Heisenberg model: $\sim\sum_{j\neq j^{\prime}}\mathcal{V}_{j,j^{\prime}}[\hat{\sigma}^{(1)}_{j}\hat{\sigma}^{(1)}_{j^{\prime}}+\hat{\sigma}^{(2)}_{j}\hat{\sigma}^{(2)}_{j^{\prime}}+\hat{\sigma}^{(3)}_{j}\hat{\sigma}^{(3)}_{j^{\prime}}$].
@@ IV.2.1 Cosine Wave
$J_{0}(z)$ takes on values (approximately) in the interval [0.4,1][-0.4,1]. If we determine the set of $s$ such that (s+3)/(s1)(s+3)/(s-1) takes on values in this interval, we can determine the set of $\hat{H}_{XXZ}$ Hamiltonians that can be engineered into $\hat{H}_{ZZ}$ using Eq. (28). Doing this, we obtain the set $s\in(-\infty,-13/7]$. Note in particular as $s\rightarrow\pm\infty$, we have $v\rightarrow 0$, which makes sense since this corresponds to $J_{\perp}\rightarrow 0$, so that $\hat{H}_{0}=\hat{H}_{ZZ}$. The restriction to $s<0$ implies that a cosine wave can produce an Ising model from the XXZ model only if there is a relative minus sign between $J_{\perp}$ and $J_{z}$.
@@
Figure 4: The solid lines plot the surfaces of solutions for $J_{z}/J_{\perp}$ obtained from Eq. (27) plotted against the pulse strength, $v$. The pulse is assumed to be a cosine shape, Eq. (28). In the top, middle, and bottom panels the dashed lines plot the truncated approximations terminating at $p_{max}=1$, $p_{max}=8$, and $p_{max}=16$, respectively. The vertical dashed lines indicate a pole in the truncated expansion. As expected, when $v\rightarrow 0$, lower order approximations are good fits to the solution surface. On the other hand, every truncated solution incorrectly predicts the existence of values of $v$ for $s>0$ (i.e., $J_{z}$ and $J_{\perp}$ having the same sign) which are pushed further and further out to $v\rightarrow\infty$. To determine all admissible values of $v$ which effectively engineer an Ising-type interaction from a given form of $H_{XXZ}$, we must keep all terms in the expansion.
@@
Appendix D briefly reviews recent quantum simulation proposals and experiments realizing XXZ models with $\mathcal{V}_{j,j^{\prime}}=|j-j^{\prime}|^{-3}$. Trapped circular Rydberg atoms can be engineered into an XXZ model where there can be a relative minus sign between $J_{\perp}$ and $J_{z}$ [NGUYEN2018]. Polar molecules, however, defining pseudospins in the simplest setup (rotational states with total angular momenta $\mathcal{N}=0,1$ and zero $z$-component of angular momentum) do not always allow a relative minus sign between $J_{\perp}$ and $J_{z}$. We therefore find that Eq. (28) does not allow a conversion from $H_{XXZ}$ to $H_{ZZ}$ for dipolar interacting molecules. But this condition can be generically met in bipartite spin model approximations to trapped polar molecules, e.g., a nearest neighbor interaction [Lee2016] $\mathcal{V}_{j,j^{\prime}}\approx\delta_{j,j^{\prime}\pm 1}$ on a chain lattice, by performing a gauge transformation on every other site on the chain: $\sigma^{(1)}_{2j}\rightarrow-\sigma^{(1)}_{2j}$ and $\sigma^{(2)}_{2j}\rightarrow-\sigma^{(2)}_{2j}$. The gauge transformation allows a relative minus sign between $J_{\perp}$ and $J_{z}$ (on a bipartite model). We therefore conclude that for the global cosine pulse, dipolar XXZ models engineered by circular Rydberg atom arrays can be driven into an Ising regime whereas only bipartite models of interacting $\mathcal{N}=0,1$ Alkali-based polar molecules can be converted from an XXZ model to an Ising model.
@@
Figure 4 uses a strong driving cosine pulse to highlight the need to retain all orders in $p$ in Eq. (21) in a broad parameter search designed to engineer models. One could in principle use the $p_{max}=1$ term, i.e., Eq. (22), to attempt to engineer the XXZ model into an Ising model, but this leads to qualitatively incorrect results in a pulse parameter sweep. In particular, the first-order solution to Eq. (27) suggests a smooth set of values for $v$ to select from that effectively engineers all possible ratios of $J_{z}$ to $J_{\perp}$ up to $J_{z}=J_{\perp}$, as seen by the top panel in Fig. 4. However, even truncating at large but finite $p_{max}$ still predicts the existence of values of $v$ which could in principle engineer specific forms of the $H_{XXZ}$, but as one keeps more and more terms in Eq. (25), these solutions are pushed out to $v\rightarrow\infty$. For large $v$, the solutions are non-physical and this can only be demonstrated by keeping all orders in $p$ in wide parameter searches over all $v$.